Find the exact value given that sin A = -4/5 with A in quadrant IV, tan B= 7/24 with B in quadrant III, and cos C= -5/13 with C in quadrant II
Cos 2A

Answer by stanbon(75887)

Find the exact value given that sin A = -4/5 with A in quadrant IV
--------
Since sin = y/4, y = -4 and r = 5
Then x = sqrt[5^2-4^2] = sqrt[9] = 3
So cos(A) = 3/5
======================================
Cos 2A = cos^2(A)-sin^2(A) = (3/5)^2 - (4/5)^2 = -7/25
=========================================================
Cheers,
Stan H.
=============